# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ 2.3 Prompting
Fine-tuning considerably reduces computational costs compared to pre-training, but it requires further model training and collections of high-quality datasets for fine-tuning, thus still consuming some computational resources and manual labor. In contrast, the “prompting” methods efficiently align general LLMs (e.g. PaLM paLM ) to the medical domain (e.g., MedPaLM medpalm ), without training any model parameters. Popular prompting methods include Zero/Few-shot prompting, Chain-of-Thought (CoT) prompting, and Prompt Tuning.
Types Tasks Datasets Data Scale Data Source
Discriminative Tasks (Sec. 3.1) Question Answering (Sec. A.1) MedQA (USMLE) medqa 61,097 multiple-choice QA pairs 51 medical textbooks
MedMCQA Medmcqa 194k multiple-choice QA pairs AIIMS and NEET PG entrance exam
MMLU hendrycks2020measuring 300 multiple-choice QA pairs Clinical Knowledge
PubMedQA PubmedQA 273,500 question-context-answer triples PubMed PubMed ; sen2008collective
BioASQ-QA krithara2023bioasq 4,721 QA pairs BioASQ tsatsaronis2015overview
EMRQA pampari2018emrqa 400k+ QA pairs i2b2
CliCR vsuster2018clicr 105k QA pairs 12k clinical case reports
COVID-QA moller2020covid 2,019 QA pairs 147 articles (CORD-19 wang2020cord )
MASH-QA zhu2020question 34,808 QA pairs WebMD website
Health-QA zhu2019hierarchical 7,517 questions and 7,355 articles 1,235 website Patient’s articles
Entity Extraction (Sec. A.2) NCBI Disease dougan2014ncbi 6,892 entity mentions 793 PubMed abstracts
JNLPBA collier2004introduction 59,963 entity mentions 2,404 abstracts (GENIA kim2003genia )
GENIA kim2003genia 96,582 entity mentions 2,000 MEDLINE abstracts
BC5CDR li2016biocreative 28,785 entity mentions 1,500 PubMed articles
BC4CHEMD krallinger2015chemdner 84,355 entity mentions 10,000 PubMed abstracts
BioRED luo2022biored 20,419 entity mentions 600 PubMed abstracts
CMeEE zhang2021cblue 24,850 entity mentions 938 files
ADE gurulingappa2012development 21,160 entity mentions 2,972 documents
2012 i2b2 sun2013evaluating 30,690 entity mentions 310 discharge summaries
2014 i2b2/UTHealth stubbs2015annotating 57,744 entity mentions 1,304 longitudinal medical records
2018 n2c2 henry20202018 83,869 entity mentions 505 discharge summariesjohnson2016mimic
Cadec karimi2015cadec 9,111 entity mentions 1,253 posts
PGR sousa2019silver 1712 abstracts 19511 entity mentions
EU-ADR van2012eu 7,011 entity mentions 100 MEDLINE abstracts
Relation Extraction (Sec. A.3) BC5CDR li2016biocreative 3,116 relations 1,500 PubMed articles
BioRED luo2022biored 6,503 relations 600 PubMed abstracts
ADE gurulingappa2012development 13,806 relations 2,972 documents
2018 n2c2 henry20202018 59,810 relations 505 discharge summaries
2010 i2b2/VA uzuner20112010 1,748 reports clinical reports
GDA wu2019renet 869,152 relations 30,192 PubMed titles and abstracts
DDI segura2013semeval 5,021 relations 1,017 biomedical texts
GAD becker2004genetic 5,937 records genetic records
2012 i2b2 sun2013evaluating 54,560 relations 310 discharge summaries
PGR sousa2019silver 4,283 relations 1,712 abstracts
EU-ADR van2012eu 2,436 relations 100 abstracts
Text Classification (Sec. A.4) ADE gurulingappa2012development 20,967 sentences 2,972 documents
2014 i2b2/UTHealth stubbs2015annotating 1,304 records longitudinal medical records
HoC baker2016automatic 5,310 sentences 1,499 PubMed abstracts
OHSUMED hersh1994ohsumed 50,216 abstracts 50,216 medical abstracts
WNUT-2020 Task 2 nguyen2020wnut 10,000 tweets 10,000 tweets
Medical Abstracts schopf2022evaluating 14,438 abstracts 28,880 medical abstracts
MIMIC-III johnson2016mimic 52,724 discharge summaries 53,423 hospital admissions
Natural Language MedNLI romanov2018lessons 14,049 sentence pairs MIMIC-III johnson2016mimic
Inference (Sec. A.5) BioNLI bastan2022bionli 40,459 sentence pairs PubMed abstracts
Semantic Textual Similarity (Sec. A.6) MedSTS wang2020medsts 174,629 sentence pairs 3M clinical notes
2019 n2c2/OHNLP wang20202019 2,054 sentence pairs 113k clinical notes and MedSTS wang2020medsts
BIOSSES souganciouglu2017biosses 100 sentence pairs Text Analysis Conference (TAC)
Information Retrieval (Sec. A.7) TREC-COVID voorhees2021trec 8,691 3-level query-document relations CORD-19 wang2020cord
NFCorpus boteva2016full 712k 3-level query-document relations NutritionFacts.org (NF) website
BioASQ (BEIR) thakur2021beir 32,916 binary query-document relations BioASQ tsatsaronis2015overview
Generative Tasks (Sec. 3.2) Text Summarization /Generation (Sec. B.1/Sec. B.3) MIMIC-CXR johnson2019mimic 128k Findings-Impression pairs radiology reports
MIMIC-III johnson2016mimic 73k Findings-Impression pairs radiology reports
PubMed PubMed ; sen2008collective 36M+ citations and abstracts biomedical literature
PMC PMC 9,407,149 papers biomedical and life sciences literature
CORD-19 wang2020cord 140k+ papers COVID-19 literature
MentSum sotudeh2022mentsum 24,119 post-TLDR summary pairs 24,119 Reddit posts
MeQSum abacha2019summarization 1,000 question-summary pairs U.S. National Library of Medicine
Text Simplification MultiCochrane joseph2023multilingual 7,755 abstract-summary pairs Cochrane Library systematic reviews
(Sec. B.2) AutoMeTS van2020automets 3,300 sentence pairs English Wikipedia kauchak2013improving
Zero/Few-shot Prompting
aims to directly give an instruction to prompt the LLM to efficiently perform a task following the given instruction. Zero-shot prompting does not provide an example. Few-shot prompting presents the LLMs with a small number of examples or task demonstrations before requiring them to perform a task. It allows the LLMs to learn from these examples or demonstrations to accurately perform the downstream task and follow the given examples to give corresponding answers GPT-3 . Therefore, few-shot prompting allows LLMs to accurately understand and respond to medical queries. For example, MedPaLM medpalm significantly improves the downstream performance by providing the general LLM, PaLM paLM , with a small number of downstream examples such as medical QA pairs.
@@
further improves the accuracy and logic of model output, compared with Zero/Few-shot Prompting. Specifically, through prompting words, CoT aims to prompt the model to generate intermediate steps or paths of reasoning when dealing with downstream (complex) problems wei2023chainofthought . Moreover, CoT can be combined with few-shot prompting by giving reasoning examples, thus enabling medical LLMs to give reasoning processes when generating responses. For tasks involving complex reasoning, such as medical QA, CoT has been shown to effectively improve model performance medpalm ; medpalm2 . Medical LLMs, such as DeID-GPT liu2023deid , MedPaLM medpalm , and MedPrompt nori2023can , use CoT prompting to assist them in simulating a diagnostic thought process, thus providing more transparent and interpretable predictions or diagnoses. In particular, MedPrompt nori2023can directly prompts a general LLM, GPT-4 gpt-4 , to outperform the fine-tuned medical LLMs on medical QA without training any model parameters.
@@
aims to improve the downstream model performance by employing both prompting and fine-tuning techniques lester2021power ; liu2021p . The prompt tuning method introduces learnable prompts, i.e. trainable continuous vectors, which can be optimized or adjusted during the fine-tuning process to better adapt to different downstream scenarios and tasks. Therefore, they provide a more flexible way of prompting LLMs than the “prompting alone” methods that use discrete and fixed prompts, as described above. In contrast to traditional fine-tuning methods that train all model parameters, prompt tuning only tunes a very small set of parameters associated with the prompts themselves, instead of extensively training the model parameters. Thus, prompt tuning effectively and accurately responds to medical problems nori2023can , with minimal incurring computational cost.
@@
Figure 3: Performance (Dataset-Metric (Task)) comparison between the GPT-3.5 turbo, GPT-4, state-of-the-art task-specific lightweight models (Fine-tuned), and human experts, on seven medical tasks across eleven datasets. All data presented in our Figures originates from published and peer-reviewed literature. Please refer to the appendix for the detailed data.
@@ 3 Medical Tasks
In this section, we will introduce two popular types of downstream tasks: generative and discriminative tasks, including ten representative downstream tasks that further build up clinical applications. Table 2 summarizes the details of widely used evaluation datasets for each downstream task. Figure 3 illustrates the performance comparisons between different LLMs. For clarity, we will only cover a general discussion of those downstream tasks. The detailed definition of the downstream task and the performance comparisons can be found in our supplementary material.
@@ 3.1 Discriminative Tasks
Discriminative tasks are for categorizing or differentiating data into specific classes or categories based on given input data. They involve making distinctions between different types of data, often to categorize, classify, or extract relevant information from structured text or unstructured text. The representative tasks (see Table 2) include Question Answering, Entity Extraction, Relation Extraction, Text Classification, Natural Language Inference, Semantic Textual Similarity, and Information Retrieval.
The typical input for discriminative tasks can be medical questions, clinical notes, medical documents, research papers, and patient EHRs. The output can be labels, categories, extracted entities, relationships, or answers to specific questions, which are often structured and categorized information derived from the input text. In existing LLMs, the discriminative tasks are widely studied, and used to make predictions and extract information from input text.
For example, based on medical knowledge, medical literature, or patient EHRs, the medical question answering (QA) task can provide precise answers to clinical questions, such as symptoms, treatment options, and drug interactions. This can help clinicians with making more efficient and accurate diagnoses medpalm ; medpalm2 ; thirunavukarasu2023large . Entity extraction can automatically identify and categorize critical information (i.e. entities) such as symptoms, medications, diseases, diagnoses, and lab results from patient EHRs, thus assisting in organizing and managing patient data. The following entity linking task aims to link the identified entities in a structured knowledge base or a standardized terminology system, e.g., SNOMED CT donnelly2006snomed , UMLS bodenreider2004unified , or ICD codes world1978international . This task is critical in clinical decision support or management systems, for better diagnosis, treatment planning, and patient care.
@@ 3.2 Generative Tasks
3.3 Performance Comparisons
Figure 3 shows that some existing general LLMs (e.g. GPT-3.5-turbo and GPT-4 gpt-4 ) have achieved strong performance on existing downstream tasks. This is most noticeable for the QA task where GPT-4 (shown in the blue line in Figure 3) consistently outperforms existing task-specific fine-tuned models, and is even comparable to human experts (shown in purple line). The QA datasets of evaluation include MedQA (USMLE) medqa , PubMedQA PubmedQA , and MedMCQA Medmcqa . To better understand the QA performance of existing medical LLMs, in Figure 4, we further demonstrate the QA performance of medical LLMs on the MedQA dataset over time in different model development types. It also clearly shows that current works, e.g., MedPrompt nori2023can , have successfully proposed several prompting methods to enable LLMs to outperform human experts.
However, on the non-QA tasks, as shown in Figure 3, the existing general LLMs perform worse than the task-specific fine-tuned models. For example, on the entity extraction task using the NCBI disease dataset dougan2014ncbi , the state-of-the-art task-specific fine-tuned model BioBERT lee2020biobert achieves an F1 score of 89.36, substantially exceeding the F1 score of 56.73 by GPT-4. We hypothesize that the reason for the strong QA capability of the current general LLMs is that the QA task is close-ended; i.e. the correct answer is already provided in multiple candidates. In contrast, most non-QA tasks are open-ended where the model has to predict the correct answer from a large pool of possible candidates, or even without any candidates provided.
Overall, the comparison proves that the current general LLMs have a strong question answering capability, however, the capability on other tasks still needs to be improved. Therefore, we advocate that the evaluation of medical LLMs should be extended to a broad range of tasks including non-QA tasks, instead of being limited mainly to medical QA tasks. Hereafter, we will discuss specific clinical applications of LLMs, followed by their challenges and future directions.

### v5

@@ 2.3 Prompting
Fine-tuning considerably reduces computational costs compared to pre-training, but it requires further model training and collections of high-quality datasets for fine-tuning, thus still consuming some computational resources and manual labor. In contrast, the “prompting” methods efficiently align general LLMs (e.g. PaLM paLM) to the medical domain (e.g., MedPaLM medpalm), without training any model parameters. Popular prompting methods include In-Context Learning (ICL), Chain-of-Thought (CoT) prompting, Prompt Tuning, and Retrieval-Augmented Generation (RAG).
In-Context Learning (ICL)
aims to directly give instructions to prompt the LLM to perform a task efficiently. In general, the ICL consists of four process: task understanding, context learning, knowledge reasoning, and answer generation. First, the model must understand the specific requirements and goals of the task. Second, the model learns to understand the contextual information related to the task with argument context. Then, use the model’s internal knowledge and reasoning capabilities to understand the patterns and logic in the example. Finally, the model generates the task-related answers. The advantage of ICL is that it does not require a large amount of labeled data for fine-tuning. Based on the type and number of input examples, ICL can be divided into three categories dong2022survey. 1) One-shot Prompting: Only one example and task description are allowed to be entered. 2) Few-shot Prompting: Allows the input of multiple instances and task descriptions. 3) Zero-shot Prompting: Only task descriptions are allowed to be entered. ICL presents the LLMs making task predictions based on contexts augmented with a few examples and task demonstrations. It allows the LLMs to learn from these examples or demonstrations to accurately perform the task and follow the given examples to give corresponding answers GPT-3. Therefore, ICL allows LLMs to accurately understand and respond to medical queries. For example, MedPaLM medpalm substantially improves the task performance by providing the general LLM, PaLM paLM, with a small number of task examples such as medical QA pairs.
@@
further improves the accuracy and logic of model output, compared with In-Context Learning. Specifically, through prompting words, CoT aims to prompt the model to generate intermediate steps or paths of reasoning when dealing with downstream (complex) problems wei2023chainofthought. Moreover, CoT can be combined with few-shot prompting by giving reasoning examples, thus enabling medical LLMs to give reasoning processes when generating responses. For tasks involving complex reasoning, such as medical QA, CoT has been shown to effectively improve model performance medpalm; medpalm2. Medical LLMs, such as DeID-GPT liu2023deid, MedPaLM medpalm, and MedPrompt nori2023can, use CoT prompting to assist them in simulating a diagnostic thought process, thus providing more transparent and interpretable predictions or diagnoses. In particular, MedPrompt nori2023can directly prompts a general LLM, GPT-4 gpt-4, to outperform the fine-tuned medical LLMs on medical QA without training any model parameters.
@@
aims to improve the model performance by employing both prompting and fine-tuning techniques lester2021power; liu2021p. The prompt tuning method introduces learnable prompts, i.e. trainable continuous vectors, which can be optimized or adjusted during the fine-tuning process to better adapt to different medical scenarios and tasks. Therefore, they provide a more flexible way of prompting LLMs than the “prompting alone” methods that use discrete and fixed prompts, as described above. In contrast to traditional fine-tuning methods that train all model parameters, prompt tuning only tunes a very small set of parameters associated with the prompts themselves, instead of extensively training the model parameters. Thus, prompt tuning effectively and accurately responds to medical problems nori2023can, with minimal incurring computational cost.
@@
Retrieval-Augmented Generation (RAG)
enhances the performance of LLMs by integrating external knowledge into the generation process. In detail, RAG can be used to minimize LLM’s hallucinations, obscure reasoning processes, and reliance on outdated information by incorporating external database knowledge gao2023retrieval. RAG consists of three main components: retrieval, augmentation, and generation. The retrieval component employs various indexing strategies and input query processing techniques to search and top-ranked relevant information from an external knowledge base. The retrieved external data is then augmented into the LLM’s prompt, providing additional context and grounding for the generated response. By directly updating the external knowledge base, RAG mitigates the risk of catastrophic forgetting associated with model weight modifications, making it particularly suitable for domains with low error tolerance and rapidly evolving information, such as the medical field. In contrast to traditional fine-tuning methods, RAG enables the timely incorporation of new medical information without compromising the model’s previously acquired knowledge, ensuring the generated outputs remain accurate and up-to-date in the face of evolving medical challenges. Most recently, researchers proposed the first benchmark MIRAGE xiong2024benchmarking based on medical information RAG, including 7,663 questions from five medical QA datasets, which has been established to both steer research and facilitate the practical deployment of medical RAG systems
In RAG, retrieval can be achieved by calculating the similarity between the embeddings of the question and document chunks, where the semantic representation capability of embedding models plays a key role. Recent research has introduced prominent embedding models such as AngIEli2023angle, Voyagewang2023voyager, and BGEbge_m3. In addition to embedding, the retrieval process can be optimized via various strategies such as adaptive retrieval, recursive retrieval, and iterative retrieval shao2023enhancing; trivedi2022interleaving; asai2023self. Several recent works have demonstrated the effectiveness of RAG in medical and pharmaceutical domains. Almanaczakka2024almanac is a large language framework augmented with retrieval capabilities for medical guidelines and treatment recommendations, surpassing the performance of ChatGPT on clinical scenario evaluations, particularly in terms of completeness and safety. Another work QA-RAG kim2024rag employs RAG with LLM for pharmaceutical regulatory tasks, where the model searches for relevant guideline documents and provides answers based on the retrieved guidelines. Chat-Orthopedistshi2023retrieval, a retrieval-augmented LLM, assists adolescent idiopathic scoliosis (AIS) patients and their families in preparing for meaningful discussions with clinicians by providing accurate and comprehensible responses to patient inquiries, leveraging AIS domain knowledge.
2.4 Discussion
This section discusses the principles of medical LLMs, including three types of methods for building models: pre-training, fine-tuning, and prompting. To meet the needs of practical medical applications, users can choose proper medical LLMs according to the magnitude of their own computing resources. Companies or institutes with massive computing resources can either pre-train an application-level medical LLM from scratch or fine-tune existing open-source general LLM models (e.g. LLaMA meta_llama3) using large-scale medical data. The results in existing literature (e.g. Med-PaLM2 medpalm2, MedAlpaca han2023medalpaca and Clinical Camel toma2023clinical) have shown that fine-tuning general LLMs on medical data can boost their performance of medical tasks. For example, Clinical Camel toma2023clinical, which is fine-tuned on the LLaMA-2-70B touvron2023llama2 model, even outperforms GPT-4 toma2023clinical. However, for small enterprises or individuals with certain computing resources, combining with the understanding of medical tasks and a reasonable combination of ICL, prompting engineering, and RAG, to prompt black-box LLMs may also achieve miraculous results. For example, MedPrompt nori2023can stimulates the commercial LLM GPT-4 gpt-4 through an appropriate combination of prompt strategies to achieve comparable or even better results than fine-tuned medical LLMs (e.g. Med-PaLM2 medpalm2) and human experts, suggesting that a mix of prompting strategies is an efficient and green solution in the medical domain rather than fine-tuning.
Figure 3: Performance (Dataset-Metric (Task)) comparison between the GPT-3.5 turbo, GPT-4, state-of-the-art task-specific lightweight models (Fine-tuned), and human experts, on seven medical tasks across eleven datasets. All data presented in our Figures originates from published and peer-reviewed literature. Please refer to the supplementary material for the detailed data.
@@ 3 Medical Tasks
In this section, we will introduce two popular types of medical machine learning tasks: generative and discriminative tasks, including ten representative tasks that further build up clinical applications. Figure 3 illustrates the performance comparisons between different LLMs. For clarity, we will only cover a general discussion of those tasks. The detailed definition of the task and the performance comparisons can be found in our supplementary material.
@@ 3.1 Discriminative Tasks
Discriminative tasks are for categorizing or differentiating data into specific classes or categories based on given input data. They involve making distinctions between different types of data, often to categorize, classify, or extract relevant information from structured text or unstructured text. The representative tasks include Question Answering, Entity Extraction, Relation Extraction, Text Classification, Natural Language Inference, Semantic Textual Similarity, and Information Retrieval.
The typical input for discriminative tasks can be medical questions, clinical notes, medical documents, research papers, and patient EHRs. The output can be labels, categories, extracted entities, relationships, or answers to specific questions, which are often structured and categorized information derived from the input text. In existing LLMs, the discriminative tasks are widely studied and used to make predictions and extract information from input text.
For example, based on medical knowledge, medical literature, or patient EHRs, the medical question answering (QA) task can provide precise answers to clinical questions, such as symptoms, treatment options, and drug interactions. This can help clinicians make more efficient and accurate diagnoses medpalm; medpalm2; thirunavukarasu2023large. Entity extraction can automatically identify and categorize critical information (i.e. entities) such as symptoms, medications, diseases, diagnoses, and lab results from patient EHRs, thus assisting in organizing and managing patient data. The following entity linking task aims to link the identified entities in a structured knowledge base or a standardized terminology system, e.g., SNOMED CT donnelly2006snomed, UMLS bodenreider2004unified, or ICD codes world1978international. This task is critical in clinical decision support or management systems, for better diagnosis, treatment planning, and patient care.
Performance Comparisons
Figure 3 shows that some existing general LLMs (e.g. GPT-3.5-turbo and GPT-4 gpt-4) have achieved strong performance on existing medical machine learning tasks. This is most noticeable for the QA task where GPT-4 (shown in the blue line in Figure 3) consistently outperforms existing task-specific fine-tuned models and is even comparable to human experts (shown in the purple line). The QA datasets of evaluation include MedQA (USMLE) medqa, PubMedQA PubmedQA, and MedMCQA Medmcqa. To better understand the QA performance of existing medical LLMs, in Figure 4, we further demonstrate the QA performance of medical LLMs on the MedQA dataset over time in different model development types. It also clearly shows that current works, e.g., MedPrompt nori2023can, have successfully proposed several prompting methods to enable LLMs to outperform human experts.
However, on the non-QA tasks, as shown in Figure 3, the existing general LLMs perform worse than the task-specific fine-tuned models. For example, on the entity extraction task using the NCBI disease dataset dougan2014ncbi, the state-of-the-art task-specific fine-tuned model BioBERT lee2020biobert achieves an F1 score of 89.36, substantially exceeding the F1 score of 56.73 by GPT-4. We hypothesize that the reason for the strong QA capability of the current general LLMs is that the QA task is close-ended; i.e. the correct answer is already provided by multiple candidates. In contrast, most non-QA tasks are open-ended where the model has to predict the correct answer from a large pool of possible candidates, or even without any candidates provided.
Overall, the comparison proves that the current general LLMs have a strong question-answering capability, however, the capability on other tasks still needs to be improved. Therefore, we advocate that the evaluation of medical LLMs should be extended to a broad range of tasks including non-QA tasks, instead of being limited mainly to medical QA tasks. Hereafter, we will discuss specific clinical applications of LLMs, followed by their challenges and future directions.
@@ 3.2 Generative Tasks
Table 3: Summary of existing medical LLMs tailored to various clinical applications, in terms of their architecture, model development, the number of parameters, the scale of PT/FT data, and the data source. M: million, B: billion. PT: Pre-training. FT: Fine-tuning. ICL: In-Context Learning. CoT: Chain-of-Thought prompting. RAG: Retrieval-Augmented Generation.
Application Model Architecture Model Development # Params Data Scale Data Source
Medical Diagnosis (Sec. 4.1) Dr. Knowsgao2023leveraging GPT-3.5 ICL 154B 5820 notes MIMIC-IIIjohnson2016mimic+IN-HOUSE gao2023leveraging
DDx PaLM-2mcduff2023towards PaLM-2 FT & ICL 340B - MultiMedQAmedpalm2+MIMIC-IIIjohnson2016mimic
Med-Geminisaab2024capabilities Gemini FT & CoT - -
MedQA-R/RSsaab2024capabilities+MultiMedQAmedpalm2
MIMIC-IIIjohnson2016mimic+MultiMedBench tanno2024consensus
Med-Flamingomoor2023med ViT+LLaMA-7B FT - 600k pairs
Multimodal Textbookmoor2023med+PMC-OAmoor2023med
VQA-RADlau2018dataset+PathVQAhe2020pathvqa
LLaVA-Medli2024llava ViT+LLaMA-7B FT - 660k pairs
PMC-15Mli2024llava+VQA-RADlau2018dataset
SLAKEliu2021slake+PathVQAhe2020pathvqa
Formatting & ICD-Coding (Sec. 4.2) PLM-ICDhuang2022plm RoBERTa FT 355M 70,539 notes MIMIC-IIsaeed2002mimic+MIMIC-IIIjohnson2016mimic
DRG-LLaMAwang2024drg LLaMA-7B FT 7B 25k pairs MIMIC-IVjohnson2019mimic
ChatICDliu2023chaticd ChatGPT ICL - 10k pairs MIMIC-IIIjohnson2016mimic
LLM-codexyang2023surpassing ChatGPT+LSTM ICL - - MIMIC-IIIjohnson2016mimic
Clinical Report Generation (Sec. 4.3) ImpressionGPTma2024iterative ChatGPT ICL & RAG 110M 184k reports MIMIC-CXRjohnson2019mimic+IU X-rayOpeni
RadAdaptvan2023radadapt T5 FT 223M, 738M 80k reports MIMIC-IIIjohnson2016mimic
ChatCADwang2023chatcad GPT-3 ICL 175B 300 reports MIMIC-CXR johnson2019mimic
MAIRA-1hyland2023maira ViT+Vicuna-7B FT 8B 337k pairs MIMIC-CXRjohnson2019mimic
RadFMwu2023towards ViT+LLaMA-13B PT& FT 14B 32M pairs MedMDwu2023towards
Medical Robotics (Sec. 4.4) SuFIAmoghani2024sufia GPT-4 ICL - 4 tasks ORBIT-Surgical yu2024orbit
UltrasoundGPTxu2024enhancing GPT-4 ICL - 522 tasks -
Robotic X-raykilleen2024take GPT-4 ICL - - -
Medical Language Translation (Sec. 4.5) Medical mT5 garcia2024medical T5 PT 738M, 3B 4.5B pairs
PubMed PubMed+EMEA tiedemann2012parallel
ClinicalTrialsNLM2022a, etc.
Apollowang2024apollo Qwen PT & FT 1.8B-7B 2.5B pairs ApolloCorporawang2024apollo
BiMediXpieri2024bimedix Mistral FT 13B 1.3M pairs BiMed1.3Mpieri2024bimedix
Biomed-sumtang2023improving BART FT 406M 27k papers BioCiteDBtang2023improving
RALLguo2024retrieval BART FT & RAG 406M 63k pairs CELLStang2023improving
Medical Education (Sec. 4.6) ChatGPTopenai2023chatgpt GPT-3.5/GPT-4 ICL - - -
Med-Geminisaab2024capabilities Gemini FT & CoT - -
MedQA-R/RS saab2024capabilities+MultiMedQAmedpalm2
MIMIC-IIIjohnson2016mimic+MultiMedBench tanno2024consensus
Mental Health Support (Sec. 4.7) PsyChatqiu2023psychat ChatGLM FT 6B 350k pairs Xinglingqiu2023psychat+Smilechatqiu2023psychat
ChatCounselorliu2023chatcounselor Vicuna FT 7B 8k instructions Psych8Kliu2023chatcounselor
Mental-LLMxu2024mental Alpaca, FLAN-T5 FT & ICL 7B, 11B 31k pairs
Dreadditturcan2019dreaddit+DepSeveritynaseem2022early+SDCNLhaque2021deep
CSSRS-Suicidegaur2019knowledge+Red-Samsampath2022data
Twt-60Usersjamil2017monitoring+SADmauriello2021sad
Medical Inquiry and Response (Sec. 4.8) AMIEtu2024towards PaLM2 FT 340B >2M pairs
MedQAmedqa+MultiMedBenchtanno2024consensus
MIMIC-IIIjohnson2016mimic+real-world diagloguetu2024towards
Healthcare Copilotren2024healthcare ChatGPT ICL - - MedDialog ren2024healthcare
Conversational Diagnosissun2024conversational GPT-4/LLaMA ICL - 40k pairs MIMIC-IVjohnson2020mimiciv